# Deep Neural Network-Based Voltage Prediction for Alkali-Metal-Ion Battery Materials

### v1

@@ Abstract
Accurately predicting the voltage of battery materials is essential for advancing energy storage technologies and designing more efficient, high-performance batteries. In this study, we developed a deep neural network (DNN) model to predict the average voltage of materials used in Li-ion, Na-ion, and other alkali-metal-ion batteries. A comprehensive dataset was compiled from the Materials Project, incorporating a diverse set of features, including structural, physical, chemical, electronic, and thermodynamic properties, along with battery-specific descriptors. These features were utilized to construct a robust DNN model aimed at facilitating the discovery of novel battery materials. The model performance evaluated through 10-fold cross validation achieving an R2 value of 0.99 and a mean absolute error (MAE) of 0.069 V on the validation dataset. This high level of accuracy underscores the model’s capability to capture the complex relationship between material properties and electrochemical performance. Furthermore, the results demonstrate that machine learning approaches, particularly DNN models, can provide rapid and reliable voltage predictions without relying on computationally expensive first-principles calculations. This research shows how systematic data analysis can accelerate the identification and improvement of next-generation battery materials. By focusing on key material properties and using computational methods, our study lays the groundwork for faster development of new energy storage solutions.
@@ 1 Introduction
In different disciplines of science and technology, the rational design of modern materials is the ultimate goal. Over the past one to two decades, tremendous effort has been made by the materials science community to compile the large dataset of materials with different properties from multiple sources1, 2, 3 so as to provide easy access of database repository to scientists and engineers to aid in novel materials discovery using artificial intelligence (AI) and machine learning (ML). In every domain such as material science, medical science, social science, advance technology, etc., there is a growing demand for new high performance, highly efficient and robust materials. To meet this demand, researchers have intensified their efforts in various fields of material science, such as high temperature superconductivity4, 5, 6, solar cell 7, 8, 9, 10, energy storage and beyond.
@@
Since last couple of decades, the Li-ion battery (LIB) is being used as a prominent renewable energy storage device and is reigning in all major electronic applications from tiny watch to large battery in EV sector and it captures most of the shares in the commercial market11. But different environment friendly renewable energy sources (solar, wind, etc.) and their utilization augments the need for higher density and highly durable energy storage materials. High manufacturing cost, low abundance of Li metal and paucity of other constituent metallic elements used in the electrode and other factors hinder further commercialization of existing battery technology to meet the current demand11. To overcome these challenges, researchers are moving towards anode-free metal-ion battery12, 13 and other earth abundant active metal-ion battery such as Na, Zn etc14, 15, 16, 17.
@@
Developing new electrode materials with high energy density and long cycle life, comparable to lithium-ion batteries (LIBs), remains a significant challenge. The process is not only technically demanding but also costly and time-consuming due to the complexities of laboratory synthesis and experimental validation. To solve these issues, researchers are trying to utilize the fast growing and high-in demand AI18 and ML19 techniques with the latest updated battery data repository such as Materials Project (MP)20, OQMD21, 22, AFlowLib23, ESP24, CMR25, NOMAD26, ICSD27, COD28, NASA29, 30, etc.
@@
Significant progress has recently been experienced in the search of new electrode materials, via prediction of battery voltages31, 32, volume change32, chemical reaction33, and formation energy33 of electrode materials using a variety of ML techniques in conjunction with the well captured clean battery database. This database3, 2 has been predominantly created using the quantum mechanical density functional theory (DFT) calculations. By combining the DFT structured database with improved and sophisticated ML algorithms, one can accelerate the discovery of modern electrode materials. Various regression and classification-based ML models have been employed for both the design and prediction of new electrode materials. For instance, Joshi et al. utilized deep neural networks (DNNs), support vector regression (SVR), and kernel ridge regression (KRR) to predict electrode voltages for metal-ion batteries, achieving a mean absolute error (MAE) of approximately 0.43 V after 10-fold cross-validation31. Similarly, Moses et al. developed a regression-based DNN model to predict the volume change of electrode materials during charge and discharge cycles, reporting an MAE of 0.47 V for average voltage predictions 32. Moreover, Louis et al. employed Graph Neural Networks (GNNs) to predict battery voltages using two different strategies, chemical reaction energy and formation energy-based approaches, obtaining MAE values in the range of 0.31 to 0.34 V33.
@@
Beyond voltage prediction, ML-driven research has also expanded towards the rational design of high-energy-density cathode and anode materials, where models are trained to screen novel compositions with optimized electrochemical stability and fast ion diffusion kinetics34. Additionally, ML techniques are being leveraged for the discovery of solid-state electrolytes (SSEs), aiming to enhance ionic conductivity while maintaining chemical and electrochemical stability 34, 35, 36. Liquid electrolytes have also been explored using ML approaches37, particularly in predicting solvation energy 38, 39, 40, electrochemical window41, 42, and ionic mobility43, 35, which are crucial for improving battery safety and performance 7, 37. Another important avenue is battery health diagnostics and lifetime prediction44, 45, 46, where ML-based models analyze degradation mechanisms, capacity fade trends, and structural stability over prolonged cycling.
As AI and ML techniques continue to evolve with larger datasets, improved feature selection strategies, and more sophisticated deep-learning architectures, their role in battery materials discovery and optimization will only grow stronger. The combination of DFT-driven data generation and ML-guided material screening has already proven to be a powerful synergy, paving the way for the next generation of high-performance, durable, and sustainable energy storage materials47.
@@
The development of high energy-density Li-ion and Na-ion battery materials has gained significant attention in both academia and industry. However, complex phase transitions during insertion and extraction, lead to challenges such as poor cycle stability, low energy density, and limited rate capability. To address these challenges, researchers have explored various strategies for optimizing Li/Na-ion cathodes. In this work, we leverage a DNN model developed using the PyTorch library to design novel layered transition metal oxide based compositions. Our model is robust and highly predictive, outperforming previously published approaches for voltage prediction33, 31, 32.
Feature engineering plays a crucial role in materials informatics, where properties derived from chemical, structural, elemental, electronic, and thermodynamic representations are used to construct feature vectors. By incorporating these effective strategies, we trained our DNN model to predict the voltage of electrode materials with high accuracy. As a result, we identified several promising cathode materials for Na-ion batteries with high voltage and energy density. To validate our predictions, we further conducted first-principles DFT simulations, confirming the potential of these materials for next-generation Na-ion batteries.
@@ 2 Methods
2.1 Collection of Data Set and their Distribution
In this work, we have collected all the metal-ion battery data from the open access Materials Project (MP, v2022.10.28) database using the Pymatgen Material Genomes tools (pymatgen)20, 48. We have extracted 4351 data for all the metal-ion batteries with DFT computed voltage and their corresponding battery features with structure. The number distribution of 10 different alkali metal-ion batteries is shown in Figure 1(a) in a bar plot and their corresponding weight percentage (pie plot) in the domain of metal-ion battery is given in Figure 1(b). It clearly shows that Li-metal ion battery dominates the dataset when compared to other metal-ion batteries. Also, we plot the average voltage distribution of the dataset (shown in Figure 1(c)), with the values ranging from 0 to 6.5 V and it signifies the number of available data distribution with certain voltage.
@@ 2.1 Collection of Data Set and their Distribution
2.2 Model Architecture and Features Generation
The features of each charge and discharge electrode of all battery materials in the MP dataset were constructed using different composition based feature generation tools such as Matminer49 and Xenonpy 50, 51. Along with key battery-related features, we also incorporate the statistical elemental properties of all constituent elements in each electrode composition. This approach helps capture the unique fingerprint of each electrode material within the dataset, enhancing the model’s ability to differentiate and predict material properties accurately. The statistical elemental feature will be different for a specific electrode material, and by considering all the statistical variables (like avg, min, max, and var) we have generated almost 262 features (see Figure 2) and used them as input for DNN model. Out of 262 features, we have generated 232 features from Xenonpy and other are extracted from Matminer and battery explorer of MP. The components of a feature vector are diverse in terms of their magnitudes, which can range from small fractions to a few thousands. Therefore, all the input features were normalized for better and more efficient training of the DNN model thereby avoiding any biased preference of a particular feature with respect to others based solely on their magnitude. The normalization was performed by effectively scaling all of the inputs to be between -1 and 1, and it was carried out on the fly while training our model and it is not done for target value. The list of statical features used to develop the DNN model is provided in the Table S1 of the Supporting Information (SI).
@@ 2.2 Model Architecture and Features Generation
2.3 Model Details
To predict the average voltage of any electrode material (cathode or anode) we implemented a DNN based on the battery electrode material properties. The architecture of the DNN was designed and implemented using PyTorch52, 53, a flexible deep learning framework. The model consisted of the following layers: input layer, hidden layers, dropout layers, and output layer. The input layer accepts a feature vector derived from material properties, where the feature size corresponds to the preprocessed dataset dimensions. The network included 5 hidden fully connected layers, each with 262 neurons. Nonlinear activation functions, specifically LeakyReLU (Leaky Rectified Linear Unit), were applied to introduce nonlinearity and also prevent overfitting. Dropout regularization was incorporated with a dropout probability of 0.2 to prevent overfitting and improve generalization by randomly dropping neurons during training. The output layer consisted of a single neuron with a linear activation function, providing the predicted value of average voltage. The Mean Squared Error (MSE) loss was used to quantify the prediction errors. MSE was chosen for its sensitivity to large errors, ensuring the model minimizes significant deviations in voltage predictions. The Adam optimizer54 was utilized for its efficiency and adaptive learning rate capabilities. The learning rate and weight decay was set to 0.0001 and 0.001.
For training, the dataset was split into training and testing subsets, with 80% of the data allocated for training and the remainder for testing. The model was trained over 1000 epochs with a batch size of 64, balancing computational efficiency and convergence. A learning rate scheduler with StepLR was employed to dynamically adjust the learning rate, ensuring better convergence. During training, both training and validation losses were monitored to ensure the model’s performance improved without overfitting. A residual plot showed that the prediction errors were well-distributed around zero, with no significant bias observed. The network architecture was defined using PyTorch’s torch.nn.Module. The model parameters were updated iteratively using backpropagation, and gradients were computed using PyTorch’s autograd functionality. After each epoch, the model was evaluated on the validation set to monitor generalization performance. The DNN demonstrated R2 value of 0.995 and 0.906 for training and testing with MSE value of 0.009 and 0.182 indicating its effectiveness in predicting the average voltage of battery materials. The model’s predictions closely matched the actual values (Figure 4), highlighting its ability to learn complex relationships between input features and the target variable.
3 Results and discussion
The performance of the ML models has been tested through different error metrics, mean absolute error (MAE) and mean square error (MSE). The MAE and MSE are defined by the following equations
@@ 3 Results and discussion
where $V_{i}^{DFT}$ represents the DFT computed voltage and $V_{i}^{ML}$ represents the machine learning predicted voltage, for the given $i^{th}$ battery sample and $N$ total number of samples in the dataset.

### v3

# Integrating Density Functional Theory with Deep Neural Networks for Accurate Voltage Prediction in Alkali-Metal-Ion Battery Materials

@@ Abstract
Accurate prediction of the voltage of battery materials plays a pivotal role in the advancement of energy storage technologies and the rational design of high-performance cathode materials. In this work, we present a deep neural network (DNN) model, built using PyTorch, to estimate the average voltage of cathode materials across Li-ion, Na-ion, and other alkali-metal-ion batteries. The model is trained on an extensive dataset from the Materials Project, incorporating a wide range of specific structural, physical, chemical, electronic, thermodynamic, and battery descriptors, ensuring a comprehensive representation of material properties. Our model exhibits strong predictive performance, as corroborated by first-principles density functional theory (DFT) calculations. The close alignment between the DNN predictions and the DFT outcomes highlights the robustness and accuracy of our machine learning framework to effectively select and identify viable battery materials. Using this validated model, we successfully proposed novel Na-ion battery compositions, with their predicted behavior confirmed by rigorous computational assessment. By seamlessly integrating data-driven prediction with first-principles validation, this study presents an effective framework that significantly accelerates the discovery and optimization of advanced battery materials, contributing to the development of more reliable and efficient energy storage technologies.
@@ 1 Introduction
In different disciplines of science and technology, the rational design of modern materials is the ultimate goal. During the past one to two decades, the materials science community has made a tremendous effort to compile the large data set of materials with different properties from multiple sources1, 2, 3 so as to provide easy access to the database repository for scientists and engineers to aid in the discovery of novel materials using artificial intelligence (AI) and machine learning (ML). In every domain such as material science, medical science, social science, advanced technology, etc., there is a growing demand for new high-performance, highly efficient, and robust materials. In recent decades, the Li-ion battery (LIB) has been used as a prominent renewable energy storage device and is dominant in all major electronic applications, from tiny watch to large battery in the EV sector and captures most of the shares in the commercial market4. However, different environmentally friendly renewable energy sources (solar, wind, etc.) and their utilization demand the need for higher density and highly durable energy storage materials. The high manufacturing cost, low abundance of Li metal and the paucity of other constituent metallic elements used in the electrode, and other factors hinder the further commercialization of existing battery technology to meet the current demand4. To overcome these challenges, researchers are moving towards anode-free metal ion batteries5, 6 and other earth abundant active metal ion batteries such as Na, Zn, etc.7, 8, 9, 10.
@@
Developing new electrode materials with high energy density and long cycle life, comparable to lithium-ion batteries (LIBs), remains a significant challenge. The process is not only technically demanding but also costly and time consuming due to the complexities of laboratory synthesis and experimental validation. To solve these issues, researchers are trying to use fast-growing and high-demand AI11 and ML12 techniques with the latest updated battery data repository such as Materials Project (MP)13, OQMD14, 15, AFlowLib16, ESP17, CMR18, NOMAD19, ICSD20, COD21, NASA22, 23, etc.
Machine learning (ML) has emerged as a powerful tool for screening potential materials (with applications in battery, catalysis, solar cells, etc.) due to its ability to significantly accelerate the discovery process compared to traditional density functional theory (DFT) calculations24. Although DFT provides high-fidelity electronic structure insights and reliable predictions of material properties, it is still computationally expensive, especially for large-scale materials exploration25. In contrast, ML models, once trained on high-quality DFT or experimental data, can predict key material properties with orders of magnitude greater efficiency, allowing for the rapid screening of vast chemical spaces24, 26. ML is particularly advantageous for identifying promising candidates early in the design process, reducing the need for exhaustive DFT calculations27. However, ML models are highly dependent on the quality and diversity of their training data, and their predictions may lack the fundamental physical interpretability of first-principles methods such as DFT28. Despite this, integrating ML with DFT in a hybrid approach can enhance both accuracy and efficiency, utilizing the robustness of DFT while leveraging the speed of ML to optimize the search for high-performance battery materials.
@@
Significant progress has recently been achieved in the search for new electrode materials, through prediction of battery voltages29, 30, volume change30, chemical reaction31, and formation energy31 of electrode materials using a variety of ML techniques in conjunction with the well-captured clean battery database. This database3, 2 has been predominantly created using quantum-mechanical DFT calculations. By combining the DFT structured database with improved and sophisticated ML algorithms, one can accelerate the discovery of modern electrode materials. Various regression- and classification-based ML models have been employed for both the design and prediction of new electrode materials. For example, Joshi et al. utilized deep neural networks (DNNs), support vector regression (SVR), and kernel ridge regression (KRR) to predict electrode voltages for metal-ion batteries, achieving a mean absolute error (MAE) of approximately 0.43 V after 10-fold cross-validation29, Mishra et al. employed XGBoost for the prediction of the average voltage in battery materials and reported a root mean square error (RMSE) of 0.41 V32. Similarly, Moses et al. developed a regression-based DNN model considering the volume change of electrode materials during the charge and discharge cycles, reporting an MAE of 0.47 V for average voltage predictions 30. Moreover, Louis et al. employed Graph Neural Networks (GNNs) to predict battery voltages using two different strategies, chemical reaction energy and formation energy-based approaches, obtaining MAE values in the range of 0.31 to 0.34 V31.
@@
Beyond voltage prediction, ML-driven research has also expanded toward the rational design of high-energy-density cathode and anode materials, where models are trained to screen novel compositions with optimized electrochemical stability and fast ion diffusion kinetics33. In addition, ML techniques are being leveraged for the discovery of solid-state electrolytes (SSEs), with the aim of improving ionic conductivity while maintaining chemical and electrochemical stability 33, 34, 35. Liquid electrolytes have also been explored using ML approaches36, particularly in predicting solvation energy 37, 38, 39, electrochemical window40, 41, and ionic mobility42, 34, which are crucial to improve battery safety and performance 43, 36. Another important avenue is battery health diagnostics and lifetime prediction44, 45, 46, where ML-based models analyze degradation mechanisms, capacity fade trends, and structural stability over prolonged cycling. As AI and ML techniques continue to evolve with larger datasets, improved feature selection strategies, and more sophisticated deep learning architectures, their role in battery material discovery and optimization will only grow stronger. The combination of DFT-driven data generation and ML-guided material screening has already proven to be a powerful synergy, paving the way for the next generation of high performance, durable, and sustainable energy storage materials47.
@@
The development of high-energy-density Li-ion and Na-ion battery materials has gained significant attention in both academia and industry. However, complex phase transitions during insertion and extraction pose challenges, such as poor cycle stability, low energy density, and limited rate capability. To address these challenges, researchers have explored various strategies for optimizing Li/Na-ion cathodes. In this work, we leverage a DNN model developed using the PyTorch library to design novel layered transition-metal oxide-based compositions. Our model is robust and highly predictive, outperforming previously published approaches for voltage prediction31, 29, 30. Feature engineering plays a crucial role in materials informatics, where properties derived from chemical, structural, elemental, electronic, and thermodynamic representations are used to construct feature vectors. By incorporating these effective strategies, we trained our DNN model to predict the voltage of electrode materials with high accuracy. As a result, we identified several promising cathode materials for Na-ion batteries with high voltage and energy density. To validate our predictions, we further conducted first-principles DFT simulations, which confirm the potential of these materials for next-generation Na-ion batteries.
2 The Rise of Na-Ion Batteries in Energy Storage
Researchers are increasingly focusing on sodium-ion batteries (Na-ion) as a promising alternative to lithium-ion (Li-ion) technology due to several key advantages in cost, resource availability, and sustainability (Figure 1). One of the most significant benefits of Na-ion batteries is their lower material cost, with a theoretical cost of \$40-77/kWh, making them 30% cheaper than LiFePO4 based lithium-ion batteries48, 49. This cost advantage is particularly crucial given the increasing costs and scarcity of lithium, which has led to an increase in production expenses for Li-ion batteries, reaching an average of \$137/kWh in 2020. In terms of energy density, Na-ion batteries offer up to 290 Wh/kg for active materials and 250-375 Wh/L volumetric energy density, making them competitive for energy storage applications. However, current practical Na-ion prototypes have a lower energy density (75-200 Wh/kg) compared to lithium-ion equivalents, but continuous research and development are expected to improve these values.
Another critical factor driving the research of Na-ion batteries is the cycle life. Although early prototypes have shown faster degradation, some advanced Na-ion battery designs have demonstrated 4,000 cycles, rivaling Li-ion batteries48, 50. This long cycle life is particularly attractive for applications such as grid storage, where energy density is less critical than longevity and cost-effectiveness. In general, the growing interest in Na-ion batteries stems from their potential to provide a sustainable, cost-effective, and scalable energy storage solution without the resource limitations associated with lithium. As research progresses, improvements in energy density, stability, and charge-discharge performance could make Na-ion batteries a viable alternative to Li-ion batteries for a wide range of applications, including renewable energy storage, electric vehicles, and large-scale power grids. As a result, the development of Na-ion battery materials has gained significant attention in both academia and industry.
Among various Na-ion cathode materials, many industries (such as HiNa, CATL, Faradian, Tiamat, etc.) have focused on layered transition metal oxides of O3 and P2-types, as they emerged as highly promising due to their superior capacity compared to polyanionic and Prussian blue-based cathodes51. However, the large ionic size of Na+ slows kinetics and induces complex phase transitions during insertion and extraction, causing challenges such as poor cycle stability, low energy density, and limited rate capability. To address these challenges, researchers have explored various strategies for optimizing Na-ion cathodes. In this work, we leverage a DNN model developed using the PyTorch library to design novel O3- and P2-type layered transition metal oxide compositions.
Figure 1: Schematic illustration highlighting the differences between Li-ion and Na-ion batteries.
@@ 3 Methods
3.1 Dataset collection and descriptors
In this work, we have collected all the data on metal-ion batteries from the Open Access Materials Project database (MP, v2022.10.28) using Pymatgen Material Genomes tools (pymatgen)13, 52. We extracted 4351 data for all the metal-ion batteries with DFT computed voltage and their corresponding battery features with structure. The distribution of the number of 10 different alkali metal-ion batteries is shown in Figure 2 (a) on a bar graph, and their corresponding weight percentage (pie graph) in the metal-ion battery domain is given in Figure 2(b). It clearly shows that the lithium metal-ion battery dominates the dataset when compared to other metal-ion batteries. In addition, we plot the average voltage distribution of the data set (shown in Figure 2 (c)), with values ranging from 0 to 6.5 V, and this indicates the number of available data distributions with certain voltage.
@@ 3.1 Dataset collection and descriptors
3.2 DNN architecture and training
The features of each charge and discharge electrode of all battery materials in the MP dataset were constructed using different composition-based feature generation tools such as Matminer53 and Xenonpy 54, 55. In addition to key battery-related features, we also incorporate the statistical elemental properties of all constituent elements in each electrode composition. This approach helps capture the unique fingerprint of each electrode material within the dataset, enhancing the model’s ability to differentiate and predict material properties accurately. The statistical elemental characteristic will be different for a specific electrode material, and by considering all statistical variables (such as avg, min, max, and var) we have generated almost 262 features (see Figure 3) and used them as input for the DNN model. Out of 262 features, we have generated 232 features from Xenonpy, and others are extracted from Matminer and battery explorer of MP. The components of a feature vector are diverse in terms of magnitude, which can range from small fractions to a few thousands. Therefore, all input features were normalized for better and more efficient training of the DNN model, thereby avoiding bias of a particular feature over others based solely on their magnitude. The normalization was performed by effectively scaling all the inputs to be between -1 and 1, and it was carried out while training our model and it is not done for the target value. The list of statistical features used to develop the DNN model is provided in Table S1 of the Supporting Information (SI).
@@ 3.2 DNN architecture and training
To predict the average voltage of any electrode material (cathode or anode), we implemented a DNN based model on the battery electrode material properties. The DNN architecture was designed and implemented using PyTorch56, 57, a flexible deep learning framework. The model consisted of the following layers: input layer, hidden layers, dropout layers, and output layer. The input layer accepts a feature vector derived from material properties, where the feature size corresponds to the dimensions of the pre-processed dataset. The network included 5 fully connected hidden layers, each with 262 neurons. Nonlinear activation functions, specifically LeakyReLU (Leaky Rectified Linear Unit), were applied to introduce nonlinearity and also prevent overfitting. Regularization of dropouts was incorporated with a dropout probability of 0.2 to prevent overfitting and improve generalization by randomly dropping neurons during training. The output layer consisted of a single neuron with a linear activation function, providing the predicted value of the average voltage. The performance of the ML models has been tested using different error metrics, mean absolute error (MAE), and mean square error (MSE). The MAE and MSE are defined by the following equations:
@@
where $V_{i}^{DFT}$ represents the voltage calculated from DFT and $V_{i}^{ML}$ represents the predicted voltage from machine learning, for the given $i^{th}$ battery sample and $N$ total number of samples in the dataset. The MSE loss was used to quantify the prediction errors. MSE was chosen for its sensitivity to large errors, ensuring that the model minimizes significant deviations in voltage predictions. The Adam optimizer58 was utilized for its efficiency and adaptive learning rate capabilities. The learning rate and weight decay were set to 0.0001 and 0.001.
For training, the data set was split into training and testing subsets, with 80% of the data allocated for training and the remainder for testing. The model was trained over 1000 epochs with a batch size of 64, balancing computational efficiency and convergence. A StepLR Learning Rate Scheduler was employed to dynamically adjust the learning rate, ensuring better convergence. During training, both training and validation losses were monitored to ensure that the model’s performance improved without overfitting. A residual plot showed that the prediction errors were well distributed around zero, with no significant bias observed. The network architecture was defined using PyTorch’s torch.nn.Module. The model parameters were updated iteratively using backpropagation, and gradients were computed using PyTorch’s autograd functionality. After each epoch, the model was evaluated on the validation set to monitor the generalization performance. The DNN demonstrated R2 value of 0.995 and 0.906 for training and testing with a MSE value of 0.009 and 0.182 indicating its effectiveness in predicting the mean voltage of battery materials. The model predictions closely matched the actual values (Figure 7), highlighting its ability to learn complex relationships between the input features and the target variable.
3.3 First-principles DFT & Boltzmann transport calculations
All DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP)59, 60 with spin-polarized ferromagnetic ordering. Formation energy calculations used the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) exchange–correlation functional61. For voltage calculations, we applied the Hubbard correction parameters (U) to account for self-interaction errors, and the DFT-D3 method62 was used to incorporate dispersion effects for all compounds of interest as the van der Waals correction is important for describing layer interactions. The U parameters used for Fe, Mn, Ni, Co, Cr and Cu were 3.9, 3.9, 6.2, 3.32, 3.7, and 5.5 eV, respectively, as obtained from Ref 63, 64. A kinetic energy cutoff of 500 eV with k-point mesh of 11 × 7 × 3 for O3 and 3 x 3 x 2 for P2-type materials was found to be sufficient to achieve convergence. During structural optimization, the atomic positions were allowed to relax until the Hellmann–Feynman forces were reduced below 0.01 eV/Å. Electronic minimization was performed with a convergence criterion of $10^{-6}$ eV. The new composition of the charge balance structure for the O3 and P2-type was built using the Supercell package 65. To generate the new compositions, here we have considered a 2 x 3 x 1 supercell lattice for O3-type materials and a 3 x 3 x 1 supercell lattice for P2-type materials.
The formation energy per formula unit (eV/f.u.) has been calculated using the formula given below 66, 32:
where $a+b+c=1$ and $M_{1}^{a}$, $M_{2}^{b}$, and $M_{3}^{c}$ represents the three different stoichiometry of metallic cations and their corresponding chemical potentials $\mu_{M_{1}}$, $\mu_{M_{2}}$, and $\mu_{M_{3}}$ respectively. Also, $\mu_{Na}$ and $\mu_{O_{2}}$ represents the chemical potential of Na bulk and O2 molecule.
The voltage (V) of the newly designed compositions was calculated using the general formula67 for multicomponent layered metal oxide which is
$V=-\frac{E_{Na_{x_{2}}MO_{2}}-E_{Na_{x_{1}}MO_{2}}-(x_{2}-x_{1})E_{Na}}{(x_{2}-x_{1})}$ (4)
where M is the multications of different stoichiometry of $M_{1}^{a}$ $M_{2}^{b}$ and $M_{3}^{c}$; $E_{Na}$ is total energy of the bulk Na and $x_{1}$ and $x_{2}$ are the different concentrations of sodium (Na) during sodiation and desodiation.
[…]
4 Results and discussion
4.1 DNN performance and cross-validation